OpenStudy (anonymous):

The period of the earth around the sun is 1 year and its distance is 150 million km from the sun. An asteroid in a circular orbit around the sun is at a distance 185 million km from the sun. Period of the asteroid is 1.369 years. What is the orbital velocity of the asteroid? Assume there are 365 days in one year. Answer in units of m/s

OpenStudy (phi):

the reasonable way to solve this is to first write 185 million km in scientific notation can you do that ?

OpenStudy (anonymous):

well I've done 1.85*10^11

OpenStudy (phi):

1.85 * 10^11 meters always keep track of the units that is the radius of a circle (the center is the sun) the total distance around is the circumference C= 2 pi r so the distance around the sun is C= 2 * 3.14 * 1.85 * 10^11 meters can you figure that out, and put the answer in scientific notation?

OpenStudy (anonymous):

Maybe but I used the \[V=2\pi r/T \] not too long ago and got an answer of 26924.1m/s

OpenStudy (phi):

yes, that is the right idea. but the first step is 2pi r what did you get for 2 pi r ?

OpenStudy (phi):

the first step in simplifying 2 * 3.14 * 1.85 * 10^11 meters is to multiply the numbers together: 2* pi *1.85 = ?

OpenStudy (phi):

you should get 2*3.14*1.85= 11.618 or 11.62 rounded now multiply by 10^11 meters 11.62*10^11 meters is the circumference we could move the decimal point of 11.62 1 to the left to get 1.162 and increase the exponent of 10^11 to 10^12 to compensate C = 1.162 * 10^12 meters

OpenStudy (phi):

speed = distance / time so speed = 1.162*10^12 meters/ 1 year (sort of like furlongs per fortnight.. not a generally used way to write speed) you want to change 1 year to seconds. How many seconds in a year ?

OpenStudy (anonymous):

31536000s

OpenStudy (phi):

we work the problem we get 60*60*24*365= 31,536,000 sec/year can you write that in scientific notation ?

OpenStudy (anonymous):

3.15*10^7

OpenStudy (phi):

ok so now we have \[ \frac{ 11.62\cdot 10^{11} \text { meters} }{3.15 \cdot 10^7 \text{ sec}}\] now simplify. 11.62/3.15 = ? 10^11 / 10^7 = ?

OpenStudy (phi):

10^11 / 10^7 = 10^{11-7} we subtract exponents when dividing = 10^4

OpenStudy (anonymous):

36888.9

OpenStudy (phi):

yes, but roughly that is 3.69 * 10^4 m/sec

OpenStudy (anonymous):

indeed